# TiM4Rec: An Efficient Sequential Recommendation Model Based on Time-Aware Structured State Space Duality Model

### v2

@@ Abstract.
Sequential recommendation represents a pivotal branch of recommendation systems, centered around dynamically analyzing the sequential dependencies between user preferences and their interactive behaviors. Despite the Transformer architecture-based models achieving commendable performance within this domain, their quadratic computational complexity relative to the sequence dimension impedes efficient modeling. In response, the innovative Mamba architecture, characterized by linear computational complexity, has emerged. Mamba4Rec further pioneers the application of Mamba in sequential recommendation. Nonetheless, Mamba 1’s hardware-aware algorithm struggles to efficiently leverage modern matrix computational units, which lead to the proposal of the improved State Space Duality (SSD), also known as Mamba 2. While the SSD4Rec successfully adapts the SSD architecture for sequential recommendation, showing promising results in high-dimensional contexts, it suffers significant performance drops in low-dimensional scenarios crucial for pure ID sequential recommendation tasks. Addressing this challenge, we propose a novel sequential recommendation backbone model, TiM4Rec, which ameliorates the low-dimensional performance loss of the SSD architecture while preserving its computational efficiency. Drawing inspiration from TiSASRec, we develop a time-aware enhancement method tailored for the linear computation demands of the SSD architecture, thereby enhancing its adaptability and achieving state-of-the-art (SOTA) performance in both low and high-dimensional modeling. The code for our model is publicly accessible at https://github.com/AlwaysFHao/TiM4Rec.
Sequential Recommendation, State Space Model(SSM), State Space Duality(SSD), Time-aware Mask Matric
@@ 1. Introduction
Sequential recommendation represents a predictive technology that leverages user’s historical interaction sequence. Unlike traditional recommendation methods—such as collaborative filtering and content based filtering—which rely heavily on static user features and historical behavior, it emphasizes capturing dynamic user behavior characteristic, which is more adept at accommodating rapid shifts in user interests and diverse consumption patterns.
@@
The landscape of sequential recommendation models has evolved significantly, transitioning from early models based on Markov chains (fpmc; hrm) to those leveraging CNNs (tang2018personalized), RNNs (li2017neural; GRU4Rec), GNNs (srgnn; chang2021sequential), notable examples include Caser (tang2018personalized) and GRU4Rec (GRU4Rec). However, these models often suffer from performance limitations inherent to their core architectures, resulting in suboptimal prediction accuracy. The introduction of the Transformer (vaswani2017attention) architecture, grounded in an attention mechanism, has marked a paradigm shift due to its superior performance and parallelizable training capabilities and made it a mainstream framework across various domains that has also significantly influenced the sequential recommendation domain. In recent years, as evidenced by models like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec) effectively adapting the success of Transformers to this area.
Despite their advancements, these early Transformer-based models did not sufficiently account for the unique data characteristics, notably the temporal shifts in user interests inherent to sequential recommendation. Subsequently, researchers introduced enhanced models with time-awareness capabilities. Models such as TiSASRec (TiSASRec) and BTMT (BTMT) extend SASRec by incorporating temporal position encoding, generating additional position encodings based on timestamps of interactions with various items, thus augmenting the expressive power of the attention score matrix. Furthermore, the TiCoSeRec (TiCoSeRec) model employs data augmentation techniques by inserting or masking items within the original interaction sequence, and its approach aims to mitigate the presence of non-uniform sub-sequences and highlight genuine interest points of users. The essence of the aforementioned methods addresses the dynamic nature of users’ interests over time. For instance, as demonstrated in Figure 1, we extract a single user’s interaction record from a real dataset, revealing significant shifts in their interests over time—an essential characteristic of user interaction behavior. Drawing inspiration from this observation, we propose a novel temporal perception enhancement method that offers substantial efficiency improvements over prior temporal enhancement methods, particularly those dominated by TiSASRec.
@@
Figure 1. User interest transfer. Over time, a noticeable shift in users’ points of interest is observed, reflecting a common characteristic of user interaction behaviors.
In contrast to the quadratic computational complexity typically associated with Transformers in the sequence dimension, the Mamba architecture offers the advantage of achieving comparable performance with linear computational complexity.With the Mamba architecture’s advantages being validated in other fields (zhu2024vision; vim-UCAS; xu2024survey; lieber2024jamba), recent studies in the realm of sequential recommendation have increasingly concentrated on applying the Mamba architecture to sequential recommendation systems (mamba4rec; bidirectionalgatedmamba4rec; echomamba4re; ssd4rec). The Mamba framework is divided into two generations: Mamba1 (gu2023mamba), which is based on SSM, and Mamba2 (SSD), which is founded on SSD. A detailed discussion of SSD and SSM is provided in Section 2.
The Mamba4Rec (mamba4rec) model is the Mamba framework’s first attempt at sequential recommendation tasks, which uses the mamba1 architecture to replace the attention mechanism module without compromising performance and significantly improving computational efficiency. Then, the SSD4Rec (ssd4rec) model is developed based on Mamba2 after Mamba4Rec. Due to the enhanced computational efficiency afforded by the Mamba2 architecture, SSD4Rec achieves superior computational performance in high-dimensional feature spaces. However, it is noteworthy that the Mamba2 architecture exhibits a performance decrement relative to Mamba1 in sequential recommendation tasks. Although SSD4Rec attains performance levels comparable to those of Mamba4Rec in high dimension settings through the implementation of variable sequences and enhancements in bidirectional SSD, since the sequential recommendation task based purely on IDs (identifiers) cannot converge efficiently in high-dimensional spaces, it performs less well than that in low-dimensional environments.
If we solely leverage the advantages of the Mamba2 architecture for high dimension modeling in sequential recommendation tasks, without considering the applicability of high-dimensional processing in pure ID-based sequential recommendation scenarios, the associated computational cost may outweigh its benefits. This is particularly pertinent given that high-dimensional modeling is relatively computationally intensive compared to its low-dimensional counterpart, and purely ID-based sequential recommendation tasks may struggle to converge efficiently in this setting. Thus, addressing the challenge of sustaining performance in low-dimensional feature spaces while leveraging the computational advantages of Mamba2 for modeling sequential recommendation tasks constitutes a critical issue necessitating further investigation.
@@
To address the above issues, this study analyzes the factors contributing to the performance degradation of SSD architecture in comparison to SSM architecture in low-dimensional sequential recommendation tasks. We introduce, for the first time, a time-aware enhancement method tailored for SSD architecture to mitigate this performance loss. This research represents the debut of Mamba architecture in the context of sequential recommendation with respect to time-aware enhancement. It is noteworthy that the time-aware enhancement method proposed in this article preserves the linear computational complexity advantage of the SSD architecture in the sequential dimension. Therefore, in comparison with the SASRec model, which represents the Transformer architecture, and the Mamba4Rec model, which represents the SSM architecture, it retains a significant advantage in terms of model training and inference speed.
@@
We designate our work as TiM4Rec and outline its primary contributions as follows:
@@
  • We conduct a pioneering exploration of a time-aware enhancement method tailored for the linear computational complexity of the Mamba architecture in the realm of sequential recommendation.
  • Through the introduction of our proposed time-aware enhancement method, we develop the TiM4Rec model, which effectively reduces performance loss in the SSD architecture relative to the SSM architecture in low-dimensional spaces, enhancing its suitability for ID modeling tasks and improving training efficiency.
  • Comprehensive experiments across three datasets demonstrate the effectiveness of the TiM4Rec model in low-dimensional spaces, showcasing superior performance compared to SSD4Rec, while maintaining the strengths of the SSD architecture in high-dimensional scenarios.
2. PRELIMINARIES
2.1. Problem Statement
Sequential recommendation methodologies leverage the user’s historical interaction sequences to forecast the subsequent potential interaction item for the user. In particular, tasks that focus solely on predicting the next item in a sequence are referred to as seq2item tasks. This work is also based on the seq2item task for its research. Let $\mathcal{U}=\left\{u_{1},u_{2},\cdots,u_{|\mathcal{U}|}\right\}$ be the set of users and $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{|\mathcal{V}|}\right\}$ be the set of items. For any user $u_{k}$, there exists an interaction sequence $\mathcal{S}_{t}^{u_{k}}=\left\{v_{1}^{u_{k}},v_{2}^{u_{k}},\cdots,v_{t}^{u_{k}}\right\}$ at time step $t$, where $v_{i}^{u_{k}}$ represents the i-th item of user $u_{k}$ interaction. Subsequently, by establishing a learnable embedding table =$\mathcal{E}=\left\{e_{1},e_{2},\cdots,e_{|\mathcal{V}|}\right\}\in\mathbb{R}^{|\mathcal{V}|\times N}$, the item IDs are converted from a sparse vector representation to a dense vector representation, where $e_{j}$ denotes the potential dense vector representation of the item $v_{j}$. We transform the input sequence $\mathcal{S}_{t}^{u_{k}}$ into the historical interaction vector representation $p^{u_{k}}_{t}$ for user $u_{k}$ through the mapping $\mathcal{F}$. Subsequently, we compute the inner product between $p^{u_{k}}_{t}$ and all item embedding vectors to obtain the predicted score ranking $r_{t}^{u_{k}}$ for the next interaction item.
$p_{t}^{u_{k}}$ $=\mathcal{F}\left(S_{t}^{u_{k}}\right)$
$r_{t}^{u_{k}}$ $=p_{t}^{u_{k}}\cdot\mathcal{E}$
Finally, we return the IDs of the top $m$ items that the user is likely to interact with, based on the predicted score ranking $r_{t}^{u_{k}}$.
@@ 2.2. State Space Models
The State Space Models(SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps the input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
@@
where $\bm{A}\in\mathbb{R}^{N\times N}$ and $\bm{B},\bm{C}\in\mathbb{R}^{N\times D}$ are learnable matrices. To enable SSM to effectively represent discrete data, it is imperative to discretize the data in accordance with the given step size $\Delta$. An effective discretization method for Equation 1 is the Zero Order Hold (ZOH) method (gu2021combining). Assuming the previous time step is $t-1$ and the current time step is $t$, Formula 1 can be solved by using the constant variation method:
@@
$=e^{\Delta\bm{A}}h(t-1)+\left(\Delta\bm{A}\right)^{-1}\left(e^{\Delta\bm{A}}-I\right)\Delta\bm{B}x\left(t\right)$
@@
The Structured State Space Model (S4) (gu2021efficiently) is developed from the vanilla SSM. By imposing a structure on the state matrix $\bm{A}\in\mathbb{R}^{D\times N}$ through HiPPO initialization (gu2020hippo), it further enhances the modeling of long-range dependencies.
Based on the foundation established by S4 model, Mamba (gu2023mamba) advances the concept of selectivity by integrating parameter matrices $\Delta\in\mathbb{R}^{T\times D}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, which vary over time. Through these methodological advancements, Mamba is capable of selectively transforming the input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ into the output sequence $\bm{Y}\in\mathbb{R}^{T\times D}$ as delineated by the following equation:
$\bm{Y}_{i}=\sum_{j}\left[\bm{C}_{i}^{\top}\left(\overline{\bm{A}}_{i}\cdot\overline{\bm{B}}_{j}\right)\right]\cdot X_{j}$
$\bm{Y}_{i}=\sum_{j}\left\{\bm{C}_{i}^{\top}\left[\left(\prod_{k=1}^{i-j}e^{\Delta_{k}^{\top}\bm{A}}\right)\cdot\left(\Delta_{j}^{\top}\bm{A}\right)^{-1}\left(e^{\Delta_{j}^{\top}\bm{A}}-\bm{I}\right)\cdot\left(\Delta_{j}^{\top}\bm{B}_{j}\right)\right]\right\}\cdot X_{j}$
Furthermore, the model enhances computational efficiency by employing parallel computation strategies facilitated through the implementation of the Hillis-Steele scan algorithm (hillis1986data). Through the aforementioned enhancements, Mamba demonstrates the capability to attain performance levels that are nearly on par with those of Transformers, while incurring only linear computational complexity with respect to the sequence dimension.

### v3

@@ Abstract.
The Sequential Recommendation modeling paradigm is shifting from Transformer to Mamba architecture, which comprises two generations: Mamba1, based on the State Space Model (SSM), and Mamba2, based on State Space Duality (SSD). Although SSD offers superior computational efficiency compared to SSM, it suffers performance degradation in sequential recommendation tasks, especially in low-dimensional scenarios that are critical for these tasks. Considering that time-aware enhancement methods are commonly employed to mitigate performance loss, our analysis reveals that the performance decline of SSD can similarly be fundamentally compensated by leveraging mechanisms in time-aware methods. Thus, we propose integrating time-awareness into the SSD framework to address these performance issues. However, integrating current time-aware methods, modeled after TiSASRec, into SSD faces the following challenges: 1) the complexity of integrating these transformer-based mechanisms with the SSD architecture, and 2) the computational inefficiency caused by the need for dimensionality expansion of time-difference modeling. To overcome these challenges, we introduce a novel Time-aware Structured Masked Matrix that efficiently incorporates time-aware capabilities into SSD. Building on this, we propose Time-Aware Mamba for Recommendation (TiM4Rec), which mitigates performance degradation in low-dimensional SSD contexts while preserving computational efficiency. This marks the inaugural application of a time-aware enhancement method specifically tailored for the Mamba architecture within the domain of sequential recommendation. Extensive experiments conducted on three real-world datasets demonstrate the superiority of our approach. The code for our model is accessible at https://github.com/AlwaysFHao/TiM4Rec.
Sequential Recommendation, State Space Model(SSM), State Space Duality(SSD), Mamba, Time-awareness
@@ 1. Introduction
Sequential Recommendation (SR) is an important branch of Recommendation Systems (RS), leveraging the user’s historical interaction sequence to predict the items that the user may interact with next (RSSurvey2005; SRSurvey2018). Unlike traditional methods—such as collaborative filtering (CollaborativeFiltering2001) and content-based filtering (VBPR2016)—that predominantly depend on static user features and past behaviors, SR focuses on capturing dynamic user behavior patterns, making it more effective in adapting to rapid shifts in user interests and diverse consumption habits (Deezer2022; SASRec2018; RSSurvey2005).
@@
The landscape of SR models has evolved significantly, transitioning from early models based on Markov chains (HRM2016; FPMC2010) to those leveraging Convolutional Neural Networks (CNNs) (TangW2018), Recurrent Neural Networks (RNNs) (GRU4Rec2016; LiRCRLM2017), Graph Neural Networks (GNNs) (ChangGZHNSJ02021; SRGNN2019), notable examples include Caser (TangW2018) and GRU4Rec (GRU4Rec2016). However, these models often suffer from performance limitations inherent to their core architectures, resulting in suboptimal prediction accuracy. The introduction of the Transformer architecture (Transformer2017), grounded in an attention mechanism, has marked a paradigm shift due to its superior performance and parallelizable training capabilities and made it a mainstream framework across various domains that have also significantly influenced the sequential recommendation domain (BTMT2024; SASRec2018; BERT4Rec2019; AFMN2022). Although Transformer-based sequential recommendation models, such as SASRec (SASRec2018), have demonstrated promising performance, their inherent quadratic computational load of the Transformer architecture poses significant challenges for efficient sequential recommendation modeling (RecBLR2024; Mamba4Rec2024; LRURec2024).
To mitigate the Transformer’s inherent computational burden, numerous models with linear complexity have been proposed (e.g., HiPPO (HiPPO2020), S4 (S4-2022), LRU (LRU2023) and RWKV (RWKV2023)). Among these, Mamba (Mamba2023) strikes a commendable balance between performance and efficiency. Mamba is derived from the State Space Model (SSM), building upon prior works (e.g., HiPPO (HiPPO2020), S4 (S4-2022), and S5 (S5-2023)), and has been widely and successfully applied across various domains (MambaNLP2024; VisionMamba2024; SSMEC2024). For efficient sequential recommendation modeling, Mamba4Rec (Mamba4Rec2024) is the first to adapt the Mamba architecture to SR. Experimental results indicate that the Mamba architecture holds greater potential than the Transformer in SR tasks (RecBLR2024; Mamba4Rec2024). Subsequently, Dao et al. (SSD2024) optimized the architecture from a parallel computing perspective, proposing a novel architecture called State Space Duality (SSD) based on Mamba, also referred to as Mamba2. SSD demonstrates superior efficiency compared to SSM and exhibits enhanced performance in high-dimensional modeling (SSD2024); however, its performance in low-dimensional settings is suboptimal (further analysis will be provided later). Although SSD4Rec (SSD4Rec2024) achieves better performance in higher dimensions than Mamba4Rec through techniques such as variable-length sequence training and bidirectional SSD, it does not outperform Mamba4Rec in lower dimensions. Pure ID sequential recommendation typically achieves optimal performance in low dimensions (BERT2019; GRU4Rec2016; SASRec2018; LRURec2024), and high-dimensional modeling contradicts the objective of efficient sequential recommendation. If we solely leverage the advantages of the Mamba2 architecture for high-dimensional modeling in SR tasks, without considering the applicability of high-dimensional processing in pure ID-based SR scenarios, the associated computational cost may outweigh its benefits. Therefore, addressing the performance deficit of the SSD architecture in low dimensions becomes a crucial challenge for efficient sequential recommendation modeling.
@@
Figure 1. User interest transfer. Over time, a noticeable shift in the user’s points of interest is observed, reflecting a common characteristic of user interaction behaviors.
User interaction behaviors typically exhibit clear temporal patterns (TiCoSeRec2023; TiSASRec2020). As illustrated in Figure 1, we extract an interaction sequence from a real-world dataset (Amazon2014), clearly demonstrating that user interests tend to change systematically over time. Previous Transformer-based SR models have not fully leveraged these temporal dynamics (MusicRec2020; TiSASRec2020; MOJITO2023). Consequently, numerous time-aware enhancement models have emerged (BTMT2024; TimelyRec2021; MEANTIME2020; TiSASRec2020; MOJITO2023; TAT4SRec2023) based on the TiSASRec (TiSASRec2020) modeling paradigm, aiming to improve the recommendation performance of Transformers by effectively incorporating time-awareness. By analyzing the reasons for the performance degradation of SSD compared to SSM, we find that, theoretically, time-aware enhancement methods can address the performance loss issue of SSD in low-dimension. Therefore, we propose integrating time-aware enhancement methods into the SSD architecture to tackle this problem. However, the existing time-aware augmentation methods encounter the following issues when integrated to SSD architectures:
Challenge I: Incompatibility with the SSD architecture. Current time-aware enhancement methods based on Transformers cannot be effectively integrated into the computational core of SSD. Specifically, time-aware enhancement methods following the TiSASRec paradigm fundamentally apply time-difference information to the attention score matrix. However, since SSD lacks an explicit representation of the attention score matrix during its computation, existing time-aware enhancement methods are incompatible with SSD’s architecture.
Challenge II: Low computational efficiency. In the paradigm of time-aware enhancement methods based on TiSASRec, it is necessary to increase the dimensionality of the time difference matrix to obtain a time-awareness score matrix, which is then integrated into the attention score matrix. This approach results in significant additional computational resource consumption, rendering existing time-aware enhancement methods ineffective for efficient sequential recommendation tasks.
@@
To address the aforementioned challenges, we introduce, for the first time, a time-aware enhancement method specifically tailored for SSD architectures that maintain computational efficiency while mitigating performance degradation issues. Specifically, we propose the Time-aware Structured Masked Matrix, which integrates time-awareness into the SSD architecture. This method involves scalar-level transformations along the sequence dimension to convert time difference information into a structured mask matrix suitable for SSD architectures. Compared to traditional time-aware enhancement approaches, our method significantly improves computational efficiency. It is noteworthy that the time-aware enhancement method proposed in this paper preserves the linear computational complexity advantage of the SSD architecture in the sequence length. Therefore, in comparison with the SASRec model (SASRec2018), which represents the Transformer architecture, and the Mamba4Rec model (Mamba4Rec2024), which represents the SSM architecture, it retains a significant advantage in terms of model training and inference speed.
@@
We designate our work as TiM4Rec(Time-aware Mamba for Recommendation) and outline its primary contributions as follows:
@@
  • To the best of our knowledge, we conduct a pioneering exploration of a time-aware enhancement method tailored for the linear computational complexity of the Mamba architecture in the field of sequential recommendation.
  • Our theoretical analysis reveals the underlying reasons for the performance degradation of SSD compared to SSM. In response, we propose a targeted time-aware enhancement method designed to mitigate the performance loss of SSD in sequential recommendation tasks.
  • We integrate time-awareness into the SSD architecture by constructing a Time-aware Structured Masked Matrix. In contrast to the time-aware enhancement method modeled on TiSASRec, which requires a computational complexity of $O(T^{2}N)$, our approach achieves efficient time-awareness with only $O(T)$ computational complexity.
  • Experiments on three real-world public datasets demonstrate the superiority of our proposed TiM4Rec compared with Transformer-based and SSM-based baselines in terms of both effectiveness and efficiency.
2. RELATED WORK
2.1. Sequential Recommendation
Previous research on sequential recommendation systems focused primarily on utilizing Recurrent Neural Networks (RNNs) (GRU4Rec2016; LiRCRLM2017). However, due to the performance limitations of RNNs and their inefficiency in parallel computation, an increasing number of successful studies in sequential recommendation have shifted towards employing Transformer-based approaches (BTMT2024; SASRec2018; BERT4Rec2019; AFMN2022). The most representative model is SASRec (SASRec2018), which captures user interest features from user interaction sequences through the causal self-attention mechanism (Transformer2017). Owing to the powerful modeling capability of self-attention, SASRec achieves strong recommendation performance. However, the Transformer architecture requires a quadratic computational load with respect to the sequence length, making it inefficient for sequence modeling (we will analyze the computational complexity in detail in Section 3.6).
In the application of recommender systems, various platforms typically require real-time recommendations, as excessive latency can detrimentally impact user experience (TencentRec2015; MaNLD2020). Consequently, the investigation of efficient sequential recommendation methods has become imperative (RecBLR2024; LRURec2024). Recently, many models with linear computational loads have emerged as alternatives to the original Transformer architecture (HiPPO2020; S4-2022; LRU2023; RWKV2023; S5-2023), among which Mamba (Mamba2023) has achieved notable success and has been promoted in various other fields (MambaNLP2024; VisionMamba2024; SSMEC2024). Mamba4Rec (Mamba4Rec2024) is the first to adapt the Mamba architecture to the sequential recommendation domain. In essence, it employs a modeling method similar to SASRec, merely replacing the self-attention mechanism module in SASRec with the Mamba1 module. Since Mamba1 uses state transitions to finely control the flow of information within sequences, rather than the fully connected modeling approach of self-attention that considers each vector’s contribution equally, it better aligns with user interaction habits (i.e., recent interactions more accurately reflect users’ next interaction tendencies). Thus, Mamba4Rec achieves better performance while maintaining efficiency.
With the introduction of the Mamba2 (SSD2024) architecture, SSD4Rec (SSD4Rec2024) has been applied to the sequential recommendation domain for the first time, further enhancing time efficiency compared to Mamba4Rec. Although SSD4Rec surpasses Mamba4Rec in high-dimensional modeling through additional processes such as variable-length sequence input and bidirectional modeling, this improvement is based on the premise that Mamba4Rec experiences a performance drop in high-dimensional settings (which we will verify in the experimental section). Prioritizing high-dimensional modeling contradicts the original intention of efficient modeling. For pure ID sequential recommendation, modeling is inherently more suitable in lower dimensions.
2.2. Time-aware methods in SR
As mentioned in the introduction, user interaction behaviors typically exhibit clear temporal patterns (TiCoSeRec2023; TiSASRec2020). Consequently, researchers have introduced time-aware enhancement methods into the field of sequential recommendation. TiSASRec (TiSASRec2020) represents the first investigation of a time-aware enhanced sequential recommendation algorithm based on the Transformer architecture. In essence, TiSASRec enhances the computation of the attention score matrix by constructing a time difference matrix and generating time difference positional encodings based on the difference values. MEANTIME (MEANTIME2020) enhances the learning of various temporal encoding patterns by utilizing multiple attention heads. TimelyRec (TimelyRec2021) models heterogeneous temporal patterns at different temporal granularities. MOJITO (MOJITO2023) captures diverse fine-grained temporal patterns simultaneously by integrating an attention mechanism with Gaussian mixtures. TAT4SRec (TAT4SRec2023) employs an encoder-decoder architecture to model timestamps and interaction items separately. Although these methods achieve improved performance compared to TiSASRec, they share the same modeling paradigm as TiSASRec and suffer from common limitations such as inefficiency and restriction to attention-based mechanisms.
The methodological paradigm of TiSASRec, while capable of enhancing the performance of Transformers in sequential recommendation through the integration of time difference matrices, incurs a significant increase in computational overhead. This is due to the necessity of embedding the time difference matrices and executing attention computations before they can be incorporated into the final attention score matrix. Such an approach contradicts the principle of efficient computation. Therefore, discovering a method to directly integrate time difference matrices into the attention score matrix is crucial for applying time-aware enhancement techniques to sequential recommendation modeling efficiently.
@@ 2.3. State Space Models
The State Space Model(SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps the input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{D}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
@@
where $\bm{A}\in\mathbb{R}^{N\times N}$ and $\bm{B},\bm{C}\in\mathbb{R}^{N\times D}$ are learnable matrices. To enable SSM to effectively represent discrete data, it is imperative to discretize the data in accordance with the given step size $\Delta$. An effective discretization method for Eq. 1 is the Zero Order Hold (ZOH) method (SSM2021). Assuming the previous time step is $t-1$ and the current time step is $t$, Eq. 1 can be solved by using the constant variation method:
@@
$=\exp(\Delta\bm{A})h(t-1)+\left(\Delta\bm{A}\right)^{-1}\left(\exp(\Delta\bm{A})-I\right)\Delta\bm{B}x\left(t\right),$
@@
The Structured State Space Model (S4) (S4-2022) is developed from the vanilla SSM. By imposing a structure on the state matrix $\bm{A}\in\mathbb{R}^{D\times N}$ through HiPPO (HiPPO2020) initialization, it further enhances the modeling of long-range dependencies.
Based on the foundation established by S4 model, Mamba (Mamba2023) advances the concept of selectivity by integrating parameter matrices $\Delta\in\mathbb{R}^{T\times D}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, which vary over time. Through these methodological advancements, Mamba is capable of selectively transforming the input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ into the output sequence $\bm{Y}\in\mathbb{R}^{T\times D}$ as delineated by the following equation:
(3) $\bm{Y}_{i}$ $=\sum_{j}\left[\bm{C}_{i}^{\top}\left(\overline{\bm{A}}_{i}\cdot\overline{\bm{B}}_{j}\right)\right]\cdot X_{j},$
$\overline{\bm{A}}_{i}$ $=\prod_{k=1}^{i-j}\exp(\Delta_{k}^{\top}\bm{A}),$
$\overline{\bm{B}}_{j}$ $=\left(\Delta_{j}^{\top}\bm{A}\right)^{-1}\left(\exp(\Delta_{j}^{\top}\bm{A})-\bm{I}\right)\cdot\left(\Delta_{j}^{\top}\bm{B}_{j}\right),$
Additionally, Mamba leverages CUDA to implement the Parallel Scan algorithm (Parallelizing2018) and draws inspiration from FlashAttention (FlashAttention2022) by employing hardware-aware parallel computation strategies to enhance computational efficiency. Through the aforementioned enhancements, Mamba demonstrates the capability to attain performance levels that are nearly on par with those of Transformers, while incurring only linear computational complexity with respect to the sequence length.